OpenStudy (anonymous):

Is there a formula for a break even point? I'm given the cost and revenue equations.

OpenStudy (anonymous):

@whpalmer4 @zepdrix @radar @mathmale @mathslover

OpenStudy (anonymous):

Profit vs units plotted with Cost vs. units will give you break-even number of units where they intersect Profit = Cost there.

OpenStudy (anonymous):

Sorry, "profit" should have been "Revenue" in my response. Profit = Revenue - Cost

OpenStudy (whpalmer4):

Right. The break even point is the point at which the profit is 0, and beyond that, positive. Typically there are startup costs which have to be covered by the initial sales before one starts turning a profit.

OpenStudy (anonymous):

I think I need help setting this up. To produce x units of a religious medal costs C(x)=18x+70. The revenue is R(x)=28x. Both cost and revenue are in dollars. Find the break-even point.

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

@douglaswinslowcooper

OpenStudy (whpalmer4):

Okay, set cost = revenue and solve for the number of units.

OpenStudy (anonymous):

Then 28x=18x+70

OpenStudy (whpalmer4):

Yes.

OpenStudy (anonymous):

10x=70 x=7

OpenStudy (anonymous):

That's the break even point then

OpenStudy (anonymous):

Now I need to find the profit for 370 units

OpenStudy (anonymous):

So Profit=28x-(18x+70)?

OpenStudy (whpalmer4):

Yes, at x=7 the cost and the revenue are equal. At x > 7, the revenue - cost > 0. For example, at x = 8, revenue is 28(8) = 224. Cost is 18(8)+70 = 214. Profit = revenue-cost = 224-214=10

OpenStudy (whpalmer4):

Yes, \[P(x) = 28x-(18x+70) = 28x-18x-70 = 10x-70\]Checking our break-even work: \[P(x) = 0 = 10x-70\]\[70=10x\]\[x=7\] All consistent, that's good!

OpenStudy (whpalmer4):

You can see how the cost function is initially larger than the revenue function until the break-even point is reached